# Context-Picker: Dynamic context selection using multi-stage reinforcement learning

### v1

@@ 3.1 Problem Formulation
We cast context picking as a single-step decision problem. For each query, a retriever first returns a candidate pool of passages $\mathcal{C}=\{(c_{1},\text{id}_{1}),(c_{2},\text{id}_{2}),\dots,(c_{N},\text{id}_{N})\},$ where $c_{j}$ is a candidate passage and $\text{id}_{j}$ is its unique identifier. Together with the query $q$ and a stage-specific instruction prompt $p_{i}$, this defines the observation $o=\langle p_{i},q,\mathcal{C}\rangle.$
The action space consists of subsets of candidate identifiers. Concretely, the policy outputs a structured response $\text{output}=\langle r,a\rangle,$ where $r$ is a rubric-guided natural-language rationale and $a=\{\text{id}_{i_{1}},\text{id}_{i_{2}},\dots,\text{id}_{i_{k}}\}\subseteq\{\text{id}_{1},\dots,\text{id}_{N}\}$ is the selected subset of IDs. The corresponding support set fed to the downstream generator is
$\mathcal{S}=\{\,c_{j}\mid(c_{j},\text{id}_{j})\in\mathcal{C},\ \text{id}_{j}\in a\,\},$
which ensures end-to-end consistency between the picker and the generator.
We constrain the action space so that $a$ is a valid, duplicate-free subset of candidate IDs; malformed or out-of-range selections are treated as invalid actions and receive format penalties in the reward. This discrete subset-based formulation matches the nature of evidence picking and serves as the MDP on which we apply GRPO-based training in the following subsections.
Given an observation–action pair $(o,a)$ with support set $\mathcal{S}$ and an offline-mined golden evidence set $\mathcal{S}_{\text{gold}}$ (Section 3.2), the stage-$i$ reward takes the abstract form
$R_{i}(o,a)=\underbrace{\mathrm{Cov}(\mathcal{S},\mathcal{S}_{\text{gold}})}_{\text{coverage}}\;-\;\underbrace{\mathrm{Redun_{i}}(\mathcal{S},\mathcal{S}_{\text{gold}})}_{\text{redundancy penalty}}\;-\;\gamma\,\mathbb{I}\bigl[\neg\text{format\_valid}(\mathcal{S})\bigr],$ (6)
where $\mathrm{Cov}$ measures how well $\mathcal{S}$ covers the golden evidence, $\mathrm{Redun}$ penalizes over-long or redundant selections in a stage-dependent manner, and the last term discourages invalid outputs. In Section 3.3 we instantiate (6) with a concrete design based on $\mathrm{Cov}$ and a normalized redundancy penalty (cf. Eq. (8)).
@@ 3.2 Data Curation
For each query–answer pair $(q,a)$, we perform retrieval using BM25 on the concatenation of the query and answer, i.e., on $[q;a]$ against the chunked document. The top-$k$ retrieved chunks constitute an initial candidate set $\mathcal{S}_{\text{cand}}$. We then run an answer-judge pipeline on $\mathcal{S}_{\text{cand}}$: a generator $\mathcal{G}$ produces a response $\hat{a}$ conditioned on $(q,\mathcal{S}_{\text{cand}})$, and an LLM-based judge $\mathcal{J}$ decides whether $\hat{a}$ semantically matches the gold answer $a$. If $\mathcal{J}$ deems $\mathcal{S}_{\text{cand}}$ insufficient (i.e., the answer is judged as incorrect), we discard this pair, since the retrieved evidence does not support a correct answer even before pruning.
@@
Considered that most long-context QA or retrieval datasets contain relatively few unique queries, we introduce lightweight query rewriting to enhance data diversity. For each original query $q$, we generate five semantically equivalent but lexically diverse reformulations $\{q_{i}^{\prime}\}_{i=1}^{5}$ using a language model. These rewrites preserve the meaning of the original query while varying in phrasing and focus, which helps improve linguistic diversity and reduces overfitting during RL training. During data partitioning, all rewrites of the same query are assigned to the same subset to prevent data leakage between training and evaluation data.
@@
Input: Document $D$; query $q$; gold answer $a$; retriever $\mathcal{R}$; encoder $f_{\text{emb}}$; generator $\mathcal{G}$; answer judge $\mathcal{J}$; top-$k$.
Output: Minimal sufficient set $\mathcal{S}_{\text{suf}}$.
$\mathcal{C}\leftarrow\text{Chunk}(D;f_{\text{emb}})$
$x\leftarrow[q;a]$
$\mathcal{S}_{\text{cand}}\leftarrow\text{RetrieveTopK}(x,\mathcal{C};\mathcal{R},k)$
$\hat{a}\leftarrow\mathcal{G}(q,\mathcal{S}_{\text{cand}})$
$r_{\text{full}}\leftarrow\mathcal{J}(q,\hat{a},a)$
if $r_{\text{full}}=0$ then
 return $\varnothing$
end if
$\mathcal{S}_{\text{suf}}\leftarrow\mathcal{S}_{\text{cand}}$
changed $\leftarrow\text{True}$
while changed do
   changed $\leftarrow\text{False}$
 foreach $c\in\mathcal{S}_{\text{suf}}$ do
    $\mathcal{S}^{\prime}\leftarrow\mathcal{S}_{\text{suf}}\setminus\{c\}$
    $\hat{a}^{\prime}\leftarrow\mathcal{G}(q,\mathcal{S}^{\prime})$
    $r^{\prime}\leftarrow\mathcal{J}(q,\hat{a}^{\prime},a)$
    if $r^{\prime}=1$ then
       $\mathcal{S}_{\text{suf}}\leftarrow\mathcal{S}^{\prime}$
         changed $\leftarrow\text{True}$
      end if
   end foreach
end while
return $\mathcal{S}_{\text{suf}}$
@@ 3.3 Multi-stage Reinforcement Learning
Using the curated training set from Section 3.2, this subsection describes the two-stage policy optimization component of Context-Picker, which corresponds to the right part of Figure 2; the detailed GRPO-based training loop is given in Algorithm 2.
In long-context question answering (LCQA), the challenges of what evidence to pick and how much evidence to pick are essentially a coupled combinatorial optimization problem. Selecting too few pieces of evidence risks missing key reasoning hops, while selecting too many introduces noise and attention dilution. Static strategies, such as fixed Top-$K$ sampling or single-stage reranking, struggle to simultaneously ensure recall sufficiency and input compactness.
As is shown in Figure 2, we decouple this problem into two training stages: a recall-oriented stage that emphasizes comprehensive evidence coverage, and a refinement-oriented stage that focuses on minimal sufficient selection.
@@
Stage I is designed to learn a high-recall picking behavior that prioritizes information completeness. In our setting, the downstream generator can answer a query correctly as long as the selected context set contains the key evidence that supports the reasoning chain. We formalize this notion via the offline-mined minimal sufficient evidence set $\mathcal{S}_{\text{gold}}$ (Section 3.2), which approximates the smallest subset that preserves answerability under an LLM-based judge.
@@
$R_{i}=\begin{cases}\mathrm{Cov}(\mathcal{S},\mathcal{S}_{\text{gold}})-\gamma\cdot\max\!\left(0,\frac{|\mathcal{S}|-|\mathcal{S}_{\text{gold}}|-\text{red}_{i}}{|\mathcal{S}_{\text{gold}}|+\text{red}_{i}}\right),&\text{if format\_valid}(\mathcal{S})\ \text{and}\ |\mathcal{S}|\leq|\mathcal{S}_{\text{gold}}|+\text{red}_{i},\\[6.0pt] 0,&\text{if format\_valid}(\mathcal{S})\ \text{and}\ |\mathcal{S}|>|\mathcal{S}_{\text{gold}}|+\text{red}_{i},\\[6.0pt] -1.0,&\text{if not format\_valid}(\mathcal{S}),\end{cases}$ (8)
where $i$ is training stage and $\mathrm{Cov}(\mathcal{S},\mathcal{S}_{\text{gold}})\;=\;\frac{|\mathcal{S}\cap\mathcal{S}_{\text{gold}}|}{|\mathcal{S}_{\text{gold}}|}$. The reward logic follows three principles:
@@
We implement the two stages as consecutive GRPO phases over the same curated dataset. In Algorithm 2, the hyperparameters $T_{1}$ and $T_{2}$ control the number of GRPO update steps spent in the recall-oriented Stage I and the refinement-oriented Stage II, respectively. In practice, we first train the picker with the Stage I reward (larger redundancy margin $\text{red}_{1}$) until the validation reward curve plateaus, and then switch to Stage II by continuing training from the Stage I checkpoint with the tighter margin $\text{red}_{2}$. We found that Context-Picker is robust to the exact split between $T_{1}$ and $T_{2}$ as long as Stage I is given enough updates to learn a high-recall policy; the resulting training dynamics for both stages are shown in Figure 3.
Input: Training set $\mathcal{D}=\{(q,\mathcal{C},\mathcal{S}_{\text{gold}})\}$; initial policy $\pi_{\theta}$; stage prompts $\{p_{1},p_{2}\}$; redundancy margins $\{\text{red}_{1},\text{red}_{2}\}$; group size $K$; iterations $\{T_{1},T_{2}\}$.
Output: Trained picker policy $\pi_{\theta}$.
Initialize reference policy $\pi_{\theta_{\mathrm{old}}}\leftarrow\pi_{\theta}$.
[2pt] for $i\in\{1,2\}$ do
 for $t=1$ to $T_{i}$ do
      Sample a mini-batch $\mathcal{B}\subset\mathcal{D}$.
      Initialize an empty set of GRPO training examples $\mathcal{G}\leftarrow\varnothing$.
    foreach $(q,\mathcal{C},\mathcal{S}_{\text{gold}})\in\mathcal{B}$ do
         Construct observation $o\leftarrow\langle p_{i},q,\mathcal{C}\rangle$.
         Sample a group of $K$ actions $\{\mathcal{S}_{1},\ldots,\mathcal{S}_{K}\}$ from $\pi_{\theta_{\mathrm{old}}}(\cdot\mid o)$.
         For each $\mathcal{S}_{j}$, compute reward $r_{j}\leftarrow R\!\bigl(\mathcal{S}_{j},o;\mathcal{S}_{\text{gold}},\text{red}_{i}\bigr)$ using Eq. (8).
         Add the group $\bigl(o,\{\mathcal{S}_{j},r_{j}\}_{j=1}^{K}\bigr)$ into $\mathcal{G}$.
     Update policy parameters $\theta$ using GRPO on $\mathcal{G}$ according to Eq. (5).
    $\pi_{\theta_{\mathrm{old}}}\leftarrow\pi_{\theta}$
return $\pi_{\theta}$
Algorithm 2 Two-stage GRPO training of Context-Picker
@@
At test time, Context-Picker runs in a single-pass retrieve–pick–generate pipeline. Given a question $q$ and a long document $D$, we first segment $D$ into semantically coherent chunks using the same chunker as in training: $\mathcal{C}\leftarrow\text{Chunk}(D;f_{\text{emb}}).$ We then build a candidate pool by retrieving the most relevant chunks to $q$, optionally truncating the pool so that the picker input fits within a budget: $\mathcal{C}_{\text{cand}}\leftarrow\text{TopSim}(q,\mathcal{C};B).$ Next, we construct the picker observation $o=\langle p_{\text{test}},q,\mathcal{C}_{\text{cand}}\rangle$ and sample the picker output from the learned policy $\pi_{\theta}$: $\{r,S\}\sim\pi_{\theta}(\cdot\mid o),$ where $r$ is a rubric-guided rationale and $S$ is the set of selected chunk identifiers. The final evidence set is obtained by filtering the candidate pool by the selected IDs, $\mathcal{C}_{\text{pick}}\leftarrow\{c_{j}\in\mathcal{C}_{\text{cand}}:\text{id}_{j}\in S\},$ and the downstream generator produces the answer conditioned on the picked evidence: $\hat{a}\leftarrow\mathcal{G}(q,\mathcal{C}_{\text{pick}}).$ For evaluation, we additionally report an LLM-as-judge score by comparing $\hat{a}$ against the reference answer (Section 4).
@@ 4.1 Experimental Setup
We evaluate Context-Picker on five knowledge-intensive QA benchmarks that require reasoning over long or multi-hop contexts: (1) LoCoMo (maharana2024evaluatinglongtermconversationalmemory), which contains extremely long multi-session conversations and tests long-term conversational memory; (2) MultiFieldQA (jiang2024longragenhancingretrievalaugmentedgeneration), a long-context QA dataset with diverse domains and relatively factoid-style questions; (3) HotpotQA (yang2018hotpotqadatasetdiverseexplainable), a classic multi-hop QA benchmark over Wikipedia; (4) 2WikiMQA (ho2020constructingmultihopqadataset), a multi-hop QA dataset requiring reasoning across two Wikipedia articles; and (5) MuSiQue (trivedi2022musiquemultihopquestionssinglehop), which decomposes multi-hop questions into compositional single-hop subquestions. For datasets in LongBench (bai2024longbenchbilingualmultitaskbenchmark) that do not come with ground-truth evidence annotations, we apply the offline evidence mining procedure in Algorithm 1 to construct training labels. Concretely, we first perform semantic chunking over each long document using text-embedding-ada-002 with a similarity threshold of $0.75$, and then mine sufficient and golden evidence sets for each $(q,a)$ pair.
Figure 3: Training dynamics of Context-Picker using GRPO. The curves visualize the average reward trajectories on training and validation sets during (Left) the Recall-Oriented Stage I and (Right) the Precision-Oriented Stage II. Both stages exhibit stable convergence and a narrow gap between training and validation rewards, indicating that the policy effectively learns to balance evidence coverage and compactness without overfitting.
Models and baselines.
Unless otherwise specified, Context-Picker is instantiated with Qwen3-8B as the picker backbone. For answer generation, we use Qwen3-32B as the generator model, and adopt GPT-4o-mini as an LLM-as-judge evaluator. Concretely, given a question $q$ and its candidate contexts, the picker selects a subset of evidence; the generator then produces an answer conditioned on $q$ and the selected evidence; finally, the judge model scores the predicted answer against the reference. We deliberately use different model families for generation and evaluation to mitigate overestimation bias when a model family evaluates its own outputs (panickssery2024llmevaluatorsrecognizefavor).
As baselines, we consider: (i) a non-retrieval LLM (Qwen3-8B) that directly consumes the raw document by concatenating $q$ with as much of the context as fits into its input window, without any retrieval or selection module; and (ii) a vanilla RAG pipeline, where a retriever returns top-$K$ passages that are directly concatenated and fed into the generator. For RAG we employ a strong dense retriever and report results for $K\in\{5,10\}$ (and $K=100$ on LoCoMo), which roughly match the average number of passages selected by Context-Picker under our token budget.
@@
Traditional metrics such as exact match (EM) and F1 are known to be brittle for free-form answers. For example, the answers “The cat is on the mat.” and “A cat rests on a mat.” convey essentially the same meaning but would receive a low EM/F1 score due to lexical differences, whereas “The cat is on the mat.” and “The dog is on the mat.” share substantial n-gram overlap while being factually incompatible. Following recent work on LLM-based evaluation (gu2025surveyllmasajudge), we thus adopt an LLM-as-judge protocol as our primary metric. Given a question $q$, a reference answer $a^{\star}$, and a predicted answer $\hat{a}$, a judge model returns a binary correctness label:
$\text{Judge}_{\text{ans}}(q,a^{\star},\hat{a})\in\{0,1\},$
based on a rubric that checks semantic equivalence to $a^{\star}$ and penalizes hallucinations or contradictions. We report the fraction of examples for which the judge predicts correctness, referred to as Judge Acc.

### v2

@@ 3.1 Problem Formulation
We cast context picking as a single-step decision problem. Given a query $q$ and a stage-specific prompt $p_{i}$, the retriever provides a candidate pool $\mathcal{C}=\{(c_{j},\text{id}_{j})\}_{j=1}^{N}$. The observation is $o=\langle p_{i},q,\mathcal{C}\rangle$. The policy outputs $\langle r,a\rangle$, where $r$ is a natural-language rationale and $a\subseteq\{\text{id}_{j}\}$ is a subset of selected IDs. This maps to the support set $\mathcal{S}=\{c_{j}\mid(c_{j},\text{id}_{j})\in\mathcal{C},\text{id}_{j}\in a\}$. Malformed or out-of-range selections are treated as invalid and penalized.
Given an offline-mined golden set $\mathcal{S}_{\text{gold}}$, the stage-$i$ reward is defined abstractly as:
$R_{i}(o,a)=\mathrm{Cov}(\mathcal{S},\mathcal{S}_{\text{gold}})-\mathrm{Redun}_{i}(\mathcal{S},\mathcal{S}_{\text{gold}})-\gamma\,\mathbb{I}[\neg\text{valid}],$ (4)
where $\mathrm{Cov}$ measures coverage, $\mathrm{Redun}_{i}$ is a stage-dependent redundancy penalty, and the last term penalizes format errors. We instantiate this in Eq. (6).
@@ 3.2 Data Curation
For each query–answer pair $(q,a)$, we perform retrieval using BM25 on the concatenation of the query and answer, i.e., on $[q;a]$ over the chunked document. The top-$k$ retrieved chunks constitute an initial candidate set $\mathcal{S}_{\text{cand}}$. We then run an answer-judge pipeline on $\mathcal{S}_{\text{cand}}$: a generator $\mathcal{G}$ produces a response $\hat{a}$ conditioned on $(q,\mathcal{S}_{\text{cand}})$, and an LLM-based judge $\mathcal{J}$ decides whether $\hat{a}$ semantically matches the gold answer $a$. If $\mathcal{J}$ deems $\mathcal{S}_{\text{cand}}$ insufficient (i.e., the answer is judged as incorrect), we discard this pair, since the retrieved evidence does not support a correct answer even before pruning.
@@
Considered that most long-context QA or retrieval datasets contain relatively few unique queries samples, we introduce lightweight query rewriting to enhance data diversity. For each original query $q$, we generate five semantically equivalent but lexically diverse reformulations $\{q_{i}^{\prime}\}_{i=1}^{5}$ using a language model. These rewrites preserve the meaning of the original query while varying in phrasing and focus, which helps improve linguistic diversity and reduces overfitting during RL training. During data partitioning, all rewrites of the same query are assigned to the same subset to prevent data leakage between training and evaluation data.
@@
0:  Document $D$; query $q$; gold answer $a$; retriever $\mathcal{R}$; encoder $f_{\text{emb}}$; generator $\mathcal{G}$; answer judge $\mathcal{J}$; top-$k$
0:  Minimal sufficient set $\mathcal{S}_{\text{suf}}$
1:  $\mathcal{C}\leftarrow\text{Chunk}(D;f_{\text{emb}})$ // semantic chunking
2:  $x\leftarrow[q;a]$
3:  $\mathcal{S}_{\text{cand}}\leftarrow\text{RetrieveTopK}(x,\mathcal{C};\mathcal{R},k)$ // initial evidence pool
4:  $\hat{a}\leftarrow\mathcal{G}(q,\mathcal{S}_{\text{cand}})$
5:  $r_{\text{full}}\leftarrow\mathcal{J}(q,\hat{a},a)$
6:  if $r_{\text{full}}=0$ then
7:   return $\emptyset$ // if evidence insufficient, drop
8:  end if
9:  $\mathcal{S}_{\text{suf}}\leftarrow\mathcal{S}_{\text{cand}}$ // start from full candidate set
10:  $\text{changed}\leftarrow\text{True}$
11:  while changed do
12:   $\text{changed}\leftarrow\text{False}$
13:   for each $c\in\mathcal{S}_{\text{suf}}$ do
14:    $\mathcal{S}^{\prime}\leftarrow\mathcal{S}_{\text{suf}}\setminus\{c\}$ // leave-one-out
15:    $\hat{a}^{\prime}\leftarrow\mathcal{G}(q,\mathcal{S}^{\prime})$
16:    $r^{\prime}\leftarrow\mathcal{J}(q,\hat{a}^{\prime},a)$
17:    if $r^{\prime}=1$ then
18:     $\mathcal{S}_{\text{suf}}\leftarrow\mathcal{S}^{\prime}$ // if $c$ is redundant, prune
19:     $\text{changed}\leftarrow\text{True}$
20:    end if
21:   end for
22:  end while
23:  return $\mathcal{S}_{\text{suf}}$
@@ 3.3 Multi-stage Reinforcement Learning
Given the curated training set from Section 3.2, this subsection describes the two-stage policy optimization component of Context-Picker, which corresponds to the right part of Figure 2. We detail the two-stage optimization process below.
In long-context question answering (LCQA), the challenges of what evidence to pick and how much evidence to pick are essentially a coupled combinatorial optimization problem. As is shown in Figure 2, we decouple this problem into two training stages: a recall-oriented stage that emphasizes comprehensive evidence coverage, and a refinement-oriented stage that focuses on minimal sufficient selection.
@@
Stage I is designed to learn a high-recall picking behavior that prioritizes information completeness. In our setting, the downstream generator can answer a query correctly as long as the selected context set contains the key evidence that supports the reasoning chain. We formalize this via the offline-mined minimal sufficient evidence set $\mathcal{S}_{\text{gold}}$ (Section 3.2), which approximates the smallest subset that preserves answerability under an LLM-based judge.
@@
Figure 3: Training dynamics of Context-Picker using GRPO. The curves visualize the average reward trajectories on training and validation sets during (Above) the Recall-Oriented Stage I and (Bottom) the Precision-Oriented Stage II. Both stages exhibit stable convergence and a narrow gap between training and validation rewards, indicating that the policy effectively learns to balance evidence coverage and compactness without overfitting.
@@
$\small R_{i}(o,a)=\begin{cases}\text{Cov}(\mathcal{S},\mathcal{S}_{g})-\gamma\dfrac{|\mathcal{S}|-|\mathcal{S}_{g}|}{L_{i}}&\text{if valid and }|\mathcal{S}|\leq L_{i},\\[6.0pt] 0&\text{if valid and }|\mathcal{S}|>L_{i},\\ -1.0&\text{if invalid},\end{cases}$ (6)
where $\mathcal{S}_{g}$ is the gold evidence set, $L_{i}=|\mathcal{S}_{g}|+\text{red}_{i}$ is the allowed length with a stage-dependent redundancy margin $\text{red}_{i}$.
The reward logic follows three principles:
@@
We implement the two stages as consecutive GRPO phases. We define hyperparameters $T_{1}$ and $T_{2}$ to control the number of update steps for the recall-oriented Stage I and the refinement-oriented Stage II, respectively. The training is sequential: we first initialize the policy $\pi_{\theta}$ and train with the Stage I reward for $T_{1}$ steps. We then update the reference policy and continue training with the Stage II reward for $T_{2}$ steps. In practice, we first train the picker with the Stage I reward (larger redundancy margin $\text{red}_{1}$) until the validation reward curve plateaus, and then switch to Stage II by continuing training from the Stage I checkpoint with the tighter margin $\text{red}_{2}$. We found that Context-Picker is robust to the exact split between $T_{1}$ and $T_{2}$ as long as Stage I is given enough updates to learn a high-recall policy; the resulting training dynamics for both stages are shown in Figure 3.
@@
At test time, Context-Picker follows a single-pass retrieve–pick–generate pipeline. Given query $q$ and document $D$, we generate chunks $\mathcal{C}$ and retrieve candidates $\mathcal{C}_{\text{cand}}\leftarrow\text{TopSim}(q,\mathcal{C};B)$ within budget $B$. We then sample the picker output $\{r,S\}\sim\pi_{\theta}(\cdot\mid\langle p_{\text{test}},q,\mathcal{C}_{\text{cand}}\rangle)$, where $S$ is the set of selected IDs. Finally, the generator produces answer $\hat{a}\leftarrow\mathcal{G}(q,\mathcal{C}_{\text{pick}})$ using evidence $\mathcal{C}_{\text{pick}}$ filtered by $S$, which is evaluated against the reference via LLM-as-judge (Section 4).
Method Setting LoCoMo MultiFieldQA HotpotQA 2WikiMQA MuSiQue
Standard Baselines
Standard RAG TopK=5 0.857 0.597 0.525 0.340
TopK=10 0.857 0.700 0.560 0.390
TopK=100 0.622
Advanced Baselines
Adaptive-k Dynamic $k$ 0.855 0.708 0.575 0.405
RankRAG TopK=10 0.655 0.875 0.732 0.665 0.465
DynamicRAG RL-Agent 0.675 0.860 0.740 0.685 0.505
Memory-R1 RL-Agent 0.695
Ours (Context-Picker)
Context-Picker (Stage 1) Recall-Oriented 0.681 0.873 0.741 0.621 0.476
Context-Picker (Stage 2) Precision-Oriented 0.706 0.825 0.747 0.702 0.522
Table 1: Main results on knowledge-intensive QA benchmarks. Reported metrics are LLM-as-a-judge Accuracy (Qwen3-32B). Best results are in bold.
Method Judge Acc (%) $\Delta$ vs. full
Context-Picker (full) 70.6
     w/o rationale 64.1 -6.5
     w/o redundancy 66.0 -4.6
     w/o Stage 1 56.5 -14.1
Table 2: Ablation study of Context-Picker on LoCoMo. $\Delta$ denotes absolute drops in Judge Acc (percentage points) compared to the full setting.
@@ 4.1 Experimental Setup
We evaluate Context-Picker on five knowledge-intensive QA benchmarks that require reasoning over long or multi-hop contexts: (1) LoCoMo Maharana et al. (2024), which contains extremely long multi-session conversations and tests long-term conversational memory; (2) MultiFieldQA Bai et al. (2024), a long-context QA dataset with diverse domains and relatively factoid-style questions; (3) HotpotQA Yang et al. (2018), a classic multi-hop QA benchmark over Wikipedia; (4) 2WikiMQA Ho et al. (2020), a multi-hop QA dataset requiring reasoning across two Wikipedia articles; and (5) MuSiQue Trivedi et al. (2022), which decomposes multi-hop questions into compositional single-hop subquestions. For datasets in LongBench Bai et al. (2024) that do not come with ground-truth evidence passages or IDs, we apply the offline evidence mining procedure in Algorithm 1 to construct training supervision. Concretely, we first perform semantic chunking over each long document using text-embedding-ada-002 with a similarity threshold of $0.75$, and then mine sufficient and golden evidence sets for each $(q,a)$ pair.
Baselines.
We compare Context-Picker against three categories of methods: (1) Standard Baselines: Standard RAG retrieving Top-$K$ ($K\in\{5,10,100\}$) chunks. (2) Heuristic/Reranking Methods: Adaptive-$k$ Taguchi et al. (2025), which dynamically selects $k$ via score thresholding, and RankRAG Yu et al. (2024) leverages an LLM-based ranker to rerank retrieved passages and select the top-k contexts for generation, thereby improving context ordering. (3) RL-based Agents: DynamicRAG Sun et al. (2025), which models the reranker as an RL agent to dynamically adjust the order and number of retrieved documents, and Memory-R1 Yan et al. (2025), which uses RL to learn memory management operations (ADD/UPDATE/DELETE/NOOP).
@@
Traditional metrics such as exact match (EM) and F1 are known to be brittle for free-form answers. For example, the answers “The cat is on the mat.” and “A cat rests on a mat.” convey essentially the same meaning but would receive a low EM/F1 score due to lexical differences, whereas “The cat is on the mat.” and “The dog is on the mat.” share substantial n-gram overlap while being factually incompatible. Following recent work on LLM-based evaluation Gu et al. (2025), we thus adopt an LLM-as-judge protocol as our primary metric. Given a question $q$ (generated by GPT-4o mini), a reference answer $a^{\star}$, and a predicted answer $\hat{a}$, a judge model returns a binary correctness label: $\text{Judge}_{\text{ans}}(q,a^{\star},\hat{a})\in\{0,1\},$ based on a rubric that checks semantic equivalence to $a^{\star}$ and penalizes hallucinations or contradictions. We report the fraction of examples for which the judge predicts correctness, referred to as Judge Acc.